OpenStudy (anonymous):

Write this rational number in decimal form. -3/5

OpenStudy (solomonzelman):

use your calculator. Put in 3÷5. Write the decimal negative.

OpenStudy (anonymous):

thanks, i did not know that

OpenStudy (solomonzelman):

yes, when you multiply a couple numbers, the result will be negative if there are odd negatives, and positive, if there are even negatives. For example \[2 \times (-3)=-6~~~~~~~~~~~and~~~~~~~~~~(-4)\times(-1)=4\]

OpenStudy (anonymous):

answer that represents the correct simplified expression of this term. d=3 2/3 - 5 1/6

OpenStudy (anonymous):

what about that @SolomonZelman

OpenStudy (solomonzelman):

Way 1. put the mixed numbers, in a form of an improper fraction, and then ADD/SUBTRACT. \(\LARGE\color{blue}{ \bf 3\frac{2 }{3}=\frac{(3 \times 3)+2 }{3} =\frac{11 }{3} }\) \(\LARGE\color{blue}{ \bf 5\frac{1 }{6}=\frac{(5 \times 6)+1 }{6} =\frac{31 }{6} }\) then subtract \(\LARGE\color{blue}{ \bf \frac{11 }{3} - \frac{31 }{6} }\) \(\LARGE\color{blue}{ \bf \frac{22 }{6} - \frac{31 }{6} }\) \(\LARGE\color{blue}{ \bf \frac{-9 }{6} }\) \(\LARGE\color{blue}{ \bf -1\frac{3 }{6} }\) \(\LARGE\color{blue}{ \bf -1\frac{1 }{2} }\)

OpenStudy (solomonzelman):

Way 2 from each of the whole number, which gives 2/3 - 2 1/6 2/3 - 13/6 4/6 - 13/6 -9/6 -1, 1/2